# Multi-Group Proportional Representation

### v1

@@ Appendix A Proofs from Section 3
Finally, using the fourth equation, we can find the optimal w:
$\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}^{*}=\widetilde{\textbf{w}}^{*}\implies\textbf{w}^{*}=(\mathbf{\Sigma}\mathbf{V}^{\intercal})^{-1}\widetilde{\textbf{w}}^{*}=\frac{\sqrt{\frac{mk}{m+k}}}{\|\mathbf{z}\|_{2}}(\mathbf{\Sigma}\mathbf{V}^{\intercal})^{-1}\mathbf{z}=\frac{\sqrt{\frac{mk}{m+k}}}{\|\mathbf{z}\|_{2}}(\mathbf{\Sigma}\mathbf{V}^{\intercal})^{-1}\mathbf{U}^{\intercal}\mathbf{\widetilde{a}}.$ (63)
The optimal w is then obtained by solving the linear system $\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}^{*}={\widetilde{\textbf{w}}}^{*}$.
Appendix B The MPR-aware retrieval algorithm
In this section, we discuss the convergence of MAPR , the Multi-group Aware Proportional Retrieval algorithm. As described in Section 4, the MPR-aware method is essentially a cutting plane method (also known as the Kelley-Cheney-Goldstein method [76, 77]) applied to the linear function $\mathbf{a}^{T}\mathbf{s}$; see also [78].
The algorithm begins by identifying the $k$-most similar items to a given query in $\mathcal{D}R$. Then, an MSE-minimizing oracle is called and returns the statistic (or group-denoting function) $c\in\cal{C}$ with the most disproportionate representation for the set of retrieved items. The function $c$ is then added as a constraint to a linear program that outputs a vector $\mathbf{a}\in[0,1]^{n}$ maximizing average similarity as in (10) subject to a single constraint $\left|\frac{1}{k}\sum{i=1}^{n}a_{i}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|\leq\rho$, where integer constraints are relaxed. The optimal $\mathbf{a}$ is rounded so only the largest $k$ entries are marked as 1, corresponding to a new set of retrieved items. The process repeats with an increasing number of constraints to the linear program and halts when the oracle does not return a function that violates the target MPR constraint $\rho$. In mathematical terms, we start by solving the problem
@@ Appendix B The MPR-aware retrieval algorithm
At each iteration, we test if the MPR constraint, less or equal to $\rho$, is satisfied. If not, this gives us a new inequality to be incorporated into the linear problem (B) that outputs a vector $\mathbf{a}\in[0,1]^{n}$ maximizing average similarity as in (10) subject to a single constraint $\left|\frac{1}{k}\sum_{i=1}^{n}a_{i}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|\leq\rho$, where integer constraints are relaxed. Therefore, the algorithm consists of iteratively adding a new constraint (cut) of the form
@@
Let $C$ be in infinite set and let $G(\mathbf{a})=\sup_{c\in\mathcal{C}}|g(\mathbf{a},c)|$ be the convex restriction of the optimization problem (10) given by the supremum, i.e., $G(\mathbf{a})\leq\rho,\forall\mathbf{a}\in\mathbb{R}^{n}$. If $c^{*}\in\arg\sup_{c\in C}|g(\mathbf{a}^{*},c)|$ and $\mathbf{c}$ denote the function $c^{*}$ evaluated at the points $x_{1},\dots,x_{n}$, i.e., $\mathbf{c}=[c_{*}(x_{1}),\dots,c_{*}(x_{m})]$, then $g^{*}:=\operatorname{sign}(g(\mathbf{a}^{*},c^{*}))\cdot\mathbf{c}\in\partial G(\mathbf{a}^{*})$, i.e., $g^{*}$ is a subgradient of $G$ at a point $\mathbf{a}^{*}$.
@@
$|g(\mathbf{a}^{*},c^{*})|+\operatorname{sign}(g(\mathbf{a}^{*},c^{*})){\mathbf{c}}^{\intercal}(\mathbf{a}-\mathbf{a}^{*})$ (69)
$=|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|+\frac{{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}}{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|}{\mathbf{c}}^{\intercal}(\mathbf{a}-\mathbf{a}^{*})$ (70)
$=|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|+\frac{{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}}{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|}-{\mathbf{c}}^{\intercal}\mathbf{a}^{*}-b_{c^{*}}+{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c^{*}}$ (71)
$=|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|-\frac{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|^{2}}{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|}+\operatorname{sign}(g(\mathbf{a}^{*},c^{*}))({\mathbf{c}}^{\intercal}\mathbf{a}+b_{c^{*}})$ (72)
$\leq|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c^{*}}|\leq G(\mathbf{a}),\forall\mathbf{a}\in\mathbb{R}^{n}$ (73)
@@
to the problem at each iteration until the solution satisfies all the inequalities. This reformulation demonstrates that the MPR-aware algorithm can be viewed as a traditional cutting plane method for minimizing a linear function over the polyhedron $P=\{a_{i}\in[0,1]\cap\sum_{i=1}^{n}a_{i}=k\}$. If, at a certain iteration, the algorithm finds an iterate that satisfies all the inequalities, it terminates since the lower bound provided by the cutting plane method matches the upper bound. In the case where the algorithm continues adding constraints, the accumulation point (guaranteed to exist due to the compactness of the set $P$) will converge to an optimal solution, as proven in [79, Proposition 4.1.2] or the main theorem in [76, Section 2].
@@
The integer program may not be as scalable as its convex relaxation. If we relax it to $a_{i}\in[0,1],$ this becomes a convex optimization problem solvable by standard methods (even for very large $\mathbf{C}$) and provides an upper bound to the original program for reasonable values of the MPR violation (and for which Proposition 9 applies). However, the optimal solution of the relaxed problem might not be feasible for the original problem, because it might have fractional values for some $a_{i}$, while the original problem requires these to be binary. There are certain special cases where the optimal solution of the relaxed problem is guaranteed to be integral (i.e., all $a_{i}$ are 0 or 1), in which case it is also an optimal solution for the original problem. This is the case, for example, when the constraint matrix is totally unimodular and the right-hand sides of the constraints are integral [82]. Still, we observe from empirical experiments that selecting the top k $a_{i}$ of the relaxed problem is a computationally efficient alternative to solving the integer program Mpr-Retriv$(k,\rho)$, since results from rounding produce a negligible difference to the upper bound (see Figures 8 and 9).
Now, we describe how the MPR constraint can be simplified for the class of linear functions. In this case, the MPR-aware retrieval algorithm becomes a scalable quadratic program.
B.1 MPR-aware Retrieval via Quadratic Programming
Proposition 4 allowed us to have a simple closed-form representation of MPR in the case of bounded-norm linear regression functions, i.e., when $\mathcal{C}^{\prime}\triangleq\{c\in\mathcal{C}\mid\frac{1}{m+n}\sum_{i=1}^{m+n}c(x_{i})^{2}=1\}$, and $\mathcal{C}=\left\{x\mapsto w^{\intercal}x\mid w\in\mathbb{R}^{d}\right\}$. In this case, MAPR  becomes
$\max_{\mathbf{a},\mathbf{y}}~{}~{}$ $\mathbf{a}^{\intercal}\mathbf{s}$ (75)
subject to $\sqrt{m+n}\left\|\mathbf{U}_{l}^{\intercal}\widetilde{\mathbf{a}}\right\|_{2}\leq\rho,$ (76)
$\mathbf{a}^{\intercal}\mathbf{1}=k,$ (77)
$a_{i}\in\{0,1\},$ (78)
Furthermore, the result from Proposition 4 allows us to explore the efficiency of our cutting plane algorithm. In Figures 4 and 6, we compare the quadratic optimization problem (Eqn. 75) to MAPR (a linear program, Algorithm 1) for linear regression and measure the MPR of our solutions with the closed form value from Proposition 4. As we can see, while the quadratic program traces a smoother curve, MAPR’s cutting plane approach well-approximates the closed form solution. In these experiments, we retrieve 50 items from CelebA, without a curation set. Finally, in Figures 5 and 7, we run the quadratic program and compare the MPR as measured with the closed form solution and linear regressor-approximated measurement of MPR. In other words, where Figures 4 and 6 measure the quality of our retrieval algorithm given a perfect (closed form) MPR oracle, Figures 5 and 7 measure the quality of our MSE oracle in comparison to the perfect oracle. We observe that the MSE oracle is a perfect approximator of the closed form for MPR.
Figure 4: Similarity vs MPR for Quadratic Program (Eqn. 75) and MAPR. MAPR well approximates the quadratic program along the Pareto frontier. Measured over a single query “A photo of a lawyer" for 50 retrieved samples on CelebA.
Figure 6: Additional comparisons of Quadratic Program (Eqn. 75) and MAPR over queries “chef", “nurse", “artist", “doctor", “teacher", “engineer", “architect", “scientist", and “programmer".
@@ B.1 MPR-aware Retrieval via Quadratic Programming
B.2 Computing MPR for Functions in an RKHS
Kernel methods are a generalized but practical, non-linear, and easy technique to implement algorithms for a plethora of important problems in machine learning [83, 84]. As MPR is closely related to the maximum mean discrepancy problem, a closed-form expression of MPR can be easily deduced by using results from [63, Lemma 6] when $\mathcal{C}$ is a reproducing kernel Hilbert space, which includes, for example, linear regression, logistic regression, Gaussian kernel ridge regression, and hard-margin SVMs.
@@ B.2 Computing MPR for Functions in an RKHS
Appendix C Description of Fair Retrieval Algorithms and Retrieval Statistics and Other Related Works
@@
  • PATHS (Srinivasan et al. [27]): It constructs a new embedding space built upon the pretrained image-text model CoCa and leveraged text-guided projections to extract an embedding that captures complex notions of diversity in people, upon which a diverse retrieval method (such as the MMR method detailed above) can be run to effectively increase the diversity of the returned samples.
Proportional Representation (PR) in political sciences. As described in the related works in Section 1, proportional representation (broadly construed) has a long history in social choice theory and political sciences [56, 57, 58, 59, 60] and has been a guiding principle in the design of political systems, aiming to ensure representation reflects underlying population preferences and demographics. It can be traced back to the late 18th century, with early proponents such as Honoré Gabriel Riqueti, Comte de Mirabeau, who discussed the idea in 1780. The 19th century saw significant contributions to the development of proportional representation, including the works of Thomas Wright Hill and Thomas Hare. Hare’s 1859 publication, “A Treatise on the Election of Representatives, Parliamentary and Municipal” [86], outlined a system of single transferable vote for proportional representation. In the late 19th and early 20th centuries, the works of Carl Andrae, Victor D’Hondt, and John Stuart Mill further explored and contributed to the discussion of proportional representation in the political science literature. See [56, 57, 58, 59, 60].
Similarly, in ML, our aim is proportional representation in data and outputs, ensuring equitable performance across groups, and mirroring representation goals in political systems. PR electoral systems [87], used by countries like Uruguay, Sweden, South Africa, and New Zealand, allocate legislative seats proportionally to votes received by parties. Challenges include e.g., gerrymandering [87] and the complexity of apportionment and seat allocation [88, 89].
@@
ClipCLIP Any VLM Ignore/Decorrelate Attributes
OURS Any Multigroup Proportional Representation over Attributes
@@
Table 2: Axis of comparison for ours and existing fair retrieval methods. The blue color indicates a positive feature of the method while the red color indicates a negative one.
Appendix D Additional Results.

### v2

@@ Appendix C Proofs from Section 3
Finally, using (45), we can find the optimal w:
$\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}^{*}=\widetilde{\textbf{w}}^{*}\implies\textbf{w}^{*}=(\mathbf{\Sigma}\mathbf{V}^{\intercal})^{-1}\widetilde{\textbf{w}}^{*}=\frac{\sqrt{\frac{mk}{m+k}}}{\|\mathbf{z}\|_{2}}(\mathbf{\Sigma}\mathbf{V}^{\intercal})^{-1}\mathbf{z}=\frac{\sqrt{\frac{mk}{m+k}}}{\|\mathbf{z}\|_{2}}(\mathbf{\Sigma}\mathbf{V}^{\intercal})^{-1}\mathbf{U}_{l}^{\intercal}\mathbf{\widetilde{a}}.$ (46)
The optimal w is then obtained by solving the linear system $\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}^{*}={\widetilde{\textbf{w}}}^{*}$. Finally, plugging the solution back in to our objective and canceling terms, we arrive at our solution. Note that we can replace $\mathbf{U}$ with $\mathbf{U}_{l}$ in our original SVD as $\mathbf{U}\in\mathbb{R}^{(m+n)\times l}$, so we can truncate the vectors in the null space of $\mathbf{U}$.
$\max_{w\in\mathbb{R}^{m+n}}~{}~{}$ $\mathbf{\widetilde{a}}^{T}\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}$
$=\mathbf{\widetilde{a}}^{T}\mathbf{U}_{l}\mathbf{\Sigma}\mathbf{V}^{\intercal}\frac{\sqrt{\frac{mk}{m+k}}}{\|\mathbf{z}\|_{2}}(\mathbf{\Sigma}\mathbf{V}^{\intercal})^{-1}\mathbf{U}_{l}^{\intercal}\mathbf{\widetilde{a}}$
$=\frac{\sqrt{\frac{mk}{m+k}}}{\|\mathbf{U}_{l}^{\intercal}\widetilde{a}\|_{2}}\mathbf{\widetilde{a}}^{T}\mathbf{U}_{l}\mathbf{U}_{l}^{\intercal}\mathbf{\widetilde{a}}$
$=\sqrt{\frac{mk}{m+k}}\|\mathbf{U}_{l}^{\intercal}\widetilde{a}\|_{2}$
Appendix D The Multi-group Optimized Retrieval Algorithm
In this section, we discuss the convergence of MOPR, the Multi-group Optimized Proportional Retrieval algorithm. As described in Section 4, the MOPR is essentially a cutting plane method (also known as the Kelley-Cheney-Goldstein method [78, 79]) applied to the linear function $\mathbf{a}^{T}\mathbf{s}$; see also [80].
The algorithm begins by identifying the $k$-most similar items to a given query in $\mathcal{D}_{R}$. Then, an MSE-minimizing oracle is called and returns the statistic (or group-denoting function) $c\in\cal{C}$ with the most disproportionate representation for the set of retrieved items. The function $c$ is then added as a constraint to a linear program that outputs a vector $\mathbf{a}\in[0,1]^{n}$ maximizing average similarity as in (9) subject to a single constraint $\left|\frac{1}{k}\sum{i=1}^{n}a_{i}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|\leq\rho$, where integer constraints are relaxed. The optimal $\mathbf{a}$ is rounded so only the largest $k$ entries are marked as 1, corresponding to a new set of retrieved items. The process repeats with an increasing number of constraints to the linear program and halts when the oracle does not return a function that violates the target MPR constraint $\rho$. In mathematical terms, we start by solving the problem
@@ Appendix D The Multi-group Optimized Retrieval Algorithm
At each iteration, we test if the MPR constraint, less or equal to $\rho$, is satisfied. If not, this gives us a new inequality to be incorporated into the linear problem (D) that outputs a vector $\mathbf{a}\in[0,1]^{n}$ maximizing average similarity as in (9) subject to a single constraint $\left|\frac{1}{k}\sum_{i=1}^{n}a_{i}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|\leq\rho$, where integer constraints are relaxed. Therefore, the algorithm consists of iteratively adding a new constraint (cut) of the form
@@
Let $\mathcal{C}$ be in infinite set and let $G(\mathbf{a})=\sup_{c\in\mathcal{C}}|g(\mathbf{a},c)|$ be the convex restriction of the optimization problem (9) given by the supremum, i.e., $G(\mathbf{a})\leq\rho,\forall\mathbf{a}\in\mathbb{R}^{n}$. If $c^{*}\in\arg\sup_{c\in\mathcal{C}}|g(\mathbf{a}^{*},c)|$ and $\mathbf{c}$ denote the function $c^{*}$ evaluated at the points $x_{1},\dots,x_{n}$, i.e., $\mathbf{c}=[c_{*}(x_{1}),\dots,c_{*}(x_{m})]$, then $g^{*}:=\operatorname{sign}(g(\mathbf{a}^{*},c^{*}))\cdot\mathbf{c}\in\partial G(\mathbf{a}^{*})$, i.e., $g^{*}$ is a subgradient of $G$ at a point $\mathbf{a}^{*}$.
@@
$|g(\mathbf{a}^{*},c^{*})|+\operatorname{sign}(g(\mathbf{a}^{*},c^{*})){\mathbf{c}}^{\intercal}(\mathbf{a}-\mathbf{a}^{*})$
$=|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|+\frac{{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}}{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|}{\mathbf{c}}^{\intercal}(\mathbf{a}-\mathbf{a}^{*})$
$=|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|+\frac{{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}}{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|}-{\mathbf{c}}^{\intercal}\mathbf{a}^{*}-b_{c^{*}}+{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c^{*}}$
$=|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|-\frac{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|^{2}}{|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c}|}+\operatorname{sign}(g(\mathbf{a}^{*},c^{*}))({\mathbf{c}}^{\intercal}\mathbf{a}+b_{c^{*}})$
$\leq|{\mathbf{c}}^{\intercal}\mathbf{a}+b_{c^{*}}|\leq G(\mathbf{a}),\forall\mathbf{a}\in\mathbb{R}^{n}$
@@
to the problem at each iteration until the solution satisfies all the inequalities. This reformulation demonstrates that the MOPR algorithm can be viewed as a traditional cutting plane method for minimizing a linear function over the polyhedron $P=\{a_{i}\in[0,1]\cap\sum_{i=1}^{n}a_{i}=k\}$. If, at a certain iteration, the algorithm finds an iterate that satisfies all the inequalities, it terminates since the lower bound provided by the cutting plane method matches the upper bound. In the case where the algorithm continues adding constraints, the accumulation point (guaranteed to exist due to the compactness of the set $P$) will converge to an optimal solution, as proven in [81, Proposition 4.1.2] or the main theorem in [78, Section 2].
@@
The integer program may not be as scalable as its convex relaxation. If we relax it to $a_{i}\in[0,1],$ this becomes a convex optimization problem solvable by standard methods (even for very large $\mathbf{C}$) and provides an upper bound to the original program for reasonable values of the MPR violation (and for which Proposition 5 applies). However, the optimal solution of the relaxed problem might not be feasible for the original problem, because it might have fractional values for some $a_{i}$, while the original problem requires these to be binary. There are certain special cases where the optimal solution of the relaxed problem is guaranteed to be integral (i.e., all $a_{i}$ are 0 or 1), in which case it is also an optimal solution for the original problem. This is the case, for example, when the constraint matrix is totally unimodular and the right-hand sides of the constraints are integral [84]. Still, we observe from empirical experiments that selecting the top k $a_{i}$ of the relaxed problem is a computationally efficient alternative to solving the integer program in Equation (9), since results from rounding produce a negligible difference to the upper bound (see Figures 2 and 3).
@@
D.1 Multi-group Optimized Retrieval via Quadratic Programming
Proposition 4 allowed us to have a simple closed-form representation of MPR in the case of bounded-norm linear regression functions, i.e., when $\mathcal{C}^{\prime}\triangleq\{c\in\mathcal{C}\mid\sqrt{\frac{mk}{m+k}}\sum_{i=1}^{m+n}c(x_{i})^{2}=1\}$, and $\mathcal{C}=\left\{x\mapsto w^{\intercal}x\mid w\in\mathbb{R}^{d}\right\}$. In this case, MOPR  becomes
$\max_{\mathbf{a},\mathbf{y}}~{}~{}$ $\mathbf{a}^{\intercal}\mathbf{s}$ (52)
subject to $\sqrt{\frac{mk}{m+k}}\left\|\mathbf{U}_{l}^{\intercal}\widetilde{\mathbf{a}}\right\|_{2}\leq\rho,$
$\mathbf{a}^{\intercal}\mathbf{1}=k,$
$a_{i}\in\{0,1\},$
This result allows us to explore the efficiency of the cutting plane algorithm in MOPR by comparing it to the solution to (52), which is a closed-form, convex quadratic program that can be easily optimized with existing solvers. Note that in these experiments, top-$k$ refers to the process of solving the relaxed linear program and then rounding to get a solution in $\{0,1\}^{n}$ by selecting the $k$ items with the highest score given by $\mathbf{a}$. This is different than the Top-$k$ described elsewhere in the paper, which denotes conducting vanilla retrieval without fairness considerations. In Figures 4 and 6, we compare the quadratic optimization problem (Eqn. 52) to MOPR (a linear program, Algorithm 1) for linear regression and measure the MPR of our solutions with the closed form value from Proposition 4. As we can see, while the quadratic program traces a smoother curve, MOPR’s cutting plane approach well-approximates the closed form solution. In these experiments, we retrieve 50 items from CelebA, without a curation set. Finally, in Figures 5 and 7, we run MOPR with the quadratic program oracle and compare our measurement of MPR with the closed form solution and linear regressor-approximated measurement of MPR. In other words, where Figures 4 and 6 measure the quality of our retrieval algorithm given a perfect (closed form) MPR oracle, Figures 5 and 7 measure the quality of our MSE estimator of MPR in comparison to the perfect oracle. We observe that the MSE oracle is a perfect approximator of the closed form for MPR, as expected.
Figure 4: Similarity vs MPR for Quadratic Program (Eqn. 52) and MOPR. MOPR well approximates the quadratic program along the Pareto frontier. Measured over a single query “A photo of a lawyer" for 50 retrieved samples on CelebA.
Figure 6: Additional comparisons of Quadratic Program (Eqn. 52) and MOPR over queries “chef", “nurse", “artist", “doctor", “teacher", “engineer", “architect", “scientist", and “programmer".
D.2 Computing MPR for Functions in an RKHS
Kernel methods are a generalized but practical, non-linear, and easy technique to implement algorithms for a plethora of important problems in machine learning [85, 86]. As MPR is closely related to the maximum mean discrepancy problem, a closed-form expression of MPR can be easily deduced by using results from [60, Lemma 6] when $\mathcal{C}$ is a reproducing kernel Hilbert space, which includes, for example, linear regression, logistic regression, Gaussian kernel ridge regression, and hard-margin SVMs.
@@ D.2 Computing MPR for Functions in an RKHS
Appendix E Description of Fair Retrieval Algorithms and Retrieval Statistics
@@
  • PATHS (Srinivasan et al. [27]): PATHS is a new embedding space built upon the pretrained image-text model CoCa and leveraged text-guided projections to extract an embedding that captures complex notions of diversity in people, upon which a diverse retrieval method (such as the MMR method detailed above, or even MOPR) can be run to effectively increase the diversity of the returned samples.
@@
CLIP-Clip Any VLM Ignore/Decorrelate Attributes
MOPR (ours) Any Multigroup Proportional Representation over Attributes
@@
Table 2: Axis of comparison for ours and existing fair retrieval methods. The blue color indicates a positive feature of the method, while the red color indicates a negative one.
Appendix F Connection to Social Choice Theory
In this section, we comment on the role of proportional representation in social sciences. As described in the related works in Section 1, proportional representation (broadly construed) has a long history in social choice theory and political sciences [88, 89, 90, 91, 92] and has been a guiding principle in the design of political systems, aiming to ensure representation reflects underlying population preferences and demographics. The foundations of social choice theory can be traced back to the late 18th century, with the seminal works of Jean-Charles de Borda and the Marquis de Condorcet, who developed fundamental voting methods and paradoxes that continue to influence modern computational approaches. During this same period, early proponents of proportional representation emerged, such as Honoré Gabriel Riqueti, Comte de Mirabeau, who discussed the idea in 1780.
The 19th century saw significant contributions to the development of proportional representation, including the works of Thomas Wright Hill, Thomas Hare, and Charles Dodgson (better known as Lewis Carroll), who made substantial contributions to voting theory and proportional representation systems. Hare’s 1859 publication, “A Treatise on the Election of Representatives, Parliamentary and Municipal” [93], outlined a system of single transferable vote for proportional representation. In the late 19th and early 20th centuries, the works of Carl Andrae, Victor D’Hondt, and John Stuart Mill further explored and contributed to the discussion of proportional representation in the political science literature. The field was revolutionized in the 20th century by Kenneth Arrow’s seminal work, whose impossibility theorem fundamentally shaped our understanding of social choice mechanisms. See [88, 89, 90, 91, 92].
Similarly, in ML, our aim is proportional representation in data and outputs, ensuring equitable performance across groups and mirroring representation goals in political systems. PR electoral systems [94], used by countries like Uruguay, Sweden, South Africa, and New Zealand, allocate legislative seats proportionally to votes received by parties. Challenges include, e.g., gerrymandering [94] and the complexity of apportionment and seat allocation [95, 96].
The emergence of computational social choice theory [97] in recent decades has built upon these classical foundations while introducing new algorithmic perspectives to address modern challenges. Researchers began studying algorithmic aspects of proportional representation, particularly focusing on the complexity of achieving different notions of proportionality in committee selection [95, 96]. The Hamilton and d’Hondt methods, initially developed for parliamentary seat allocation, have been extended and generalized to handle more complex scenarios. As mentioned at the end of Section 1, a notable contribution in this direction is the work of Lang and Skowron [37], who introduced multi-attribute proportional representation. Their framework generalizes classical apportionment methods to scenarios where candidates possess multiple attributes (such as gender, profession, and age), and representation goals must be satisfied across all these dimensions simultaneously. Their work provides both theoretical guarantees through approximation algorithms and practical methods for achieving proportional representation in modern committee selection problems, bridging the gap between classical political science approaches and contemporary computational challenges. In particular, they introduced the problem of multi-attribute proportional representation in committee selection. Their framework can be viewed as solving an $(m,t)$-filling problem: given $m$ buckets (representing intersectional groups) with target capacities $\{t_{1},...,t_{m}\}$ (derived from marginal probabilities), the goal is to optimally fill these buckets with appropriately labeled balls (committee members).
While groundbreaking in extending classical apportionment methods to multiple attributes, their approach differs from ours in several key aspects. First, their work considers multiple attributes, but they primarily focus on matching target proportions derived from marginal distributions for each attribute independently rather than truly addressing intersectional representation. That is, their approach aims to satisfy target proportions for each individual attribute (like gender or profession separately) but does not explicitly handle the complex interactions between attributes (like representation of specific gender-profession combinations). This is in contrast to our framework, where the function class $\mathcal{C}$ can directly capture and measure such intersectional relationships. Moreover, their formulation focuses solely on matching target distributions without additional objectives, whereas our work explicitly balances representation with retrieval utility through cosine similarity maximization. Second, while they work with pre-specified target probabilities, we estimate the desired representation from a reference dataset, making our approach more flexible and data-driven. Third, and perhaps most importantly, our MPR metric measures representation through the lens of computationally identifiable groups defined by the aforementioned function class $\mathcal{C}$ rather than explicitly enumerating all possible intersectional groups. Through $\mathcal{C}$, our framework provides a unified way to handle both simple and complex representations: it can capture discrete indicators for each attribute-value pair (like the Hamilton method; see Definition 5 in [37]) but also complex nonlinear relationships (like d’Hondt method; see Definition 6 in [37]), and goes beyond both to measure intersectional representation via decision trees, linear functions, neural networks, or more generally, any function in a reproducing kernel Hilbert space (RKHS). This allows us to handle more complex notions of representation while maintaining computational tractability. In essence, while [37] provides fundamental insights for multi-attribute committee selection, our framework offers a more general approach suitable for modern retrieval systems where both representation and relevance matter.
Appendix G Additional Results
In this section, we include additional experimental results. First, we include a comparison of runtimes for each method in Table 3. We find that MOPR provides up to a 100x speedup when compared with the most competitive retrieval algorithm, MMR, while remaining competitive to methods such as CLIP-Clip and DebiasCLIP that do not modify the retrieval algorithm at all and just use embedding similarity search. We additionally provide the similarity score for the minimum MPR achieved by each algorithm, which was used to report the results in Tables 1 and 5. These similarity scores can be found in Table 4. Even when significantly closing the MPR gap and achieving more representation than other methods, MOPR also preserves high cosine similarity with respect to the query.